Equivalent Units of Production and Related Costs

The charges to Work in Process—Assembly Department for a period, together with information concerning production, are as follows. All direct materials are placed in process at the beginning of production.

Work in Process-Assembly Department
Bal., 4,000 units, 50% completed 14,600 To Finished Goods, 92,000 units ?
Direct materials, 94,000 units @ $1.90 178,600
Direct labor 257,700
Factory overhead 100,260
Bal. ? units, 70% completed ?

Determine the following:

a. The number of units in work in process inventory at the end of the period.
units

b. Equivalent units of production for direct materials and conversion. If an amount is zero or a blank, enter in "0".

Work in Process-Assembly Department
Equivalent Units of Production for Direct Materials and Conversion Costs
Whole Units Equivalent Units Direct Materials Equivalent Units Conversion
Inventory in process, beginning
Started and completed
Transferred to finished goods
Inventory in process, ending
Total units

c. Costs per equivalent unit for direct materials and conversion. If required, round your answers to the nearest cent.

Costs Per Equivalent Unit
Direct Materials $
Conversion $

d. Cost of the units started and completed during the period.
$
Answer #1

a) Ending WIP unit = 4000+94000-92000 = 6000 Units

b) Equivalent unit

Work in Process-Assembly Department
Equivalent Units of Production for Direct Materials and Conversion Costs
Whole Units Equivalent Units Direct Materials Equivalent Units Conversion
Inventory in process, beginning 4000 0 4000*50% = 2000
Started and completed 88000 88000 88000
Transferred to finished goods 92000 88000 90000
Inventory in process, ending 6000 6000 6000*70% = 4200
Total units 98000 94000 94200

c. Costs per equivalent unit for direct materials and conversion. If required, round your answers to the nearest cent.

Direct material = 178600/94000 = 1.90

Conversion = (257700+100260)/94200 = 3.80

d) Cost of unit started and completed = (1.9+3.8)*88000 = 501600